# What Layers When: Learning to Skip Compute in LLMs with Residual Gates

### v1

@@ 4.6 End-to-End Efficiency Gains in Real-World Scenarios
Table 10 shows end-to-end latency and throughput measurements for our Llama-1b GateSkip model evaluated on GSM8K and CommonsenseQA-Gen. The results show that GateSkips theoretical FLOP savings translate into analogous real-world efficiency gains. More extensive GPU-level optimizations will likely increase these gains significantly, as our implementation does not resort to, e.g., custom GPU kernels or accelerated inference systems such as vLLM.
@@ 5 Limitations
Our study targets 1–8 B‑parameter decoder‑only LLMs and evaluates on English reasoning, translation, and general language modeling. Because of limited GPU resources, we report single‑seed results and approximate variance by 100 k‑bootstrap resampling; seed‑level sensitivity could be higher in noisier domains. We report theoretical FLOP reductions in the main text, as they more faithfully capture methodological differences and enable fair comparison across approaches, while also presenting end-to-end efficiency gains in § 4.6. An ethics statement and LLM usage is disclosed in Appendix D.
@@ C.1 Visualization Set‑up
where $H$ is the hidden dimension. We create heatmaps showing the average gate value for each token and layer, separated into Attention and MLP modules for better clarity of the patterns distinct to each module type. The resulting heatmaps are shown in Figures 2 and 4. Darker colors correspond to higher compute allocation.
@@ C.2 BOS Tokens and Punctuation as Structural Anchors
C.3 Gate Values as an Interpretability and Safety Tool
@@
C.4 Analysis of Gate Value Distribution
@@
Appendix D Ethics Statement
@@
Appendix E GateSkip Algorithms
@@
Appendix F Compute Resources
@@
Appendix G Parameter and memory overhead of GateSkip
@@
Table 13: Parameter and memory overhead of gating variants.
Appendix H Dataset
H.1 Dataset Meta Data
@@
H.2 Exact Template Used During Training
@@
Appendix I Instructions for Code Reproducibility and Access to Code
@@ I.1 Running Experiments
I.2 Collecting and Visualizing Results

### v3

@@ 4.6 End-to-End Efficiency Gains in Real-World Scenarios
Table 10 shows end-to-end latency and throughput measurements for our Llama-1b GateSkip model evaluated on GSM8K and CommonsenseQA-Gen using an optimized vLLM environment. The results show that GateSkips theoretical FLOP savings translate into analogous real-world efficiency gains. Analogously, Figure 7 in Appendix D shows consistently decreasing wall-clock time with lower FLOPs per token.
@@ 5 Limitations
Our study targets 1–8 B‑parameter decoder‑only LLMs and evaluates on English reasoning, translation, and general language modeling. We report theoretical FLOP reductions in the main text, as they more faithfully capture methodological differences and enable fair comparison across approaches, while also presenting end-to-end efficiency gains in § 4.6. An ethics statement and LLM usage is disclosed in Appendix E.
@@ C.1 Visualization Set‑up
where $H$ is the hidden dimension. We create heatmaps showing the average gate value for each token and layer, separated into Attention and MLP modules for better clarity of the patterns distinct to each module type. The resulting heatmaps are shown in Figures 2 and 5. Darker colors correspond to higher compute allocation.
@@ C.2 BOS Tokens and Punctuation as Structural Anchors
C.3 Gate Values per Token Type and Layer
To quantitatively compare how gate values differ among different token categories and layers throughout an entire corpus, we record per-layer gate values across an entire evaluation run for default GateSkip-1b and group them by token type. Specifically, as most of the model’s generation lies within its Chain-of-Thought (CoT), we record prepositions, numbers, verbs, and nouns within its CoT and juxtapose this with non-CoT tokens. Figure 4 shows heatmaps with the token type on the y-axis and layer index on the x-axis, again differentiated into Attention and MLP layers.
For attention layers, the first two and the middle layers exhibit high gate scores relative to the rest of the network. Especially layer 0 shows significantly higher importance than all other layers, while layer 13 possesses the lowest importance scores. Notably, layers six to nine have consistently high importance, perhaps corresponding to more abstract semantic operations explored in former literature (vig-belinkov-2019-analyzing; geva-etal-2021-transformer). Numbers have notably higher importance than other tokens in layers six to nine, while prepositions consistently exhibit lower importance than other token types.
Contrastingly, MLP layers generally have uniformly high gate values, indicating that all of them are needed for nearly every token. This fits the hypothesis that MLP layers encode knowledge and perform static checks for memorized facts (geva-etal-2021-transformer).
Figure 4: Average gate scores per layer and token type. The left plot corresponds to attention layers while the right one shows MLP layers.
C.4 Gate Values as an Interpretability and Safety Tool
@@
C.5 Analysis of Gate Value Distribution
@@
Appendix D FLOPs vs. Wall-Clock Time
Figure 7: FLOPs/token vs. wall-clock time/token.
Appendix E Ethics Statement
@@
Appendix F GateSkip Algorithms
@@
Appendix G Compute Resources
@@
Appendix H Parameter and memory overhead of GateSkip
Table 13: Parameter and memory overhead of gating variants.
@@ Appendix H Parameter and memory overhead of GateSkip
Appendix I Dataset
I.1 Dataset Meta Data
@@
I.2 Exact Template Used During Training
@@
Appendix J Instructions for Code Reproducibility and Access to Code
@@
J.2 Collecting and Visualizing Results